Project

General

Profile

11/06/2013 Minutes

Attendance

Mat, Michael, Joao, Gavin, Dominick, Kanika

Operations

FarDet is turned off. They are fixing hardware that did not cool correctly.
Peter is at AR doing shakedown of the DAQ.
Partition 2, APDs are installed on DCMs 3-5, 3-4 and 3-6. They will be
NDOS turned off today. DCM 1-2 taken off NDOS and becomes a teststand for new FEBs for underground.
~300 FEBs remain on NDOS.

Production

  • OnMon
    NDOS 24hr plots are not filled - they are blank. NDOS numi cana plots.
    130 events in the last week on the NDOS weekly plot. We will monitor this.
    322 events in the monthly

Michael added to the nearline pages. If partition is unknown, not set he dumps to fake partition (4) for the pages.
It's not linked but you can change p1 -> p4 to see if the data is there - Expert toggle!

  • Watchdog plots show that the raw-to-root step is now successfully completing for all runs.
    Still remain some missing raw files. Reconstruction still failing.

Outstanding Issues

  • Networking
    Appears to be resolved. Things are running again.
    OnMon is back up and running
  • Nearline plots for Time Spill Server
  • Files failed reco step in processing. Why??
  • Watchdog plots
    Question for Dominick. Seems to be an inconsistency between the latest file modification time and the plots.
    Answer is that it was just based on the final CAF files. Dominick will query the DB, find runs that happened in the last week.

Action Item: We will spend a chunk of tomorrow's session flipping through EventDisplay's for Michael's Slicer tuning study.